OpenStudy (anonymous):

The circumference of the Earth is approximately 25,000 miles. Find the diameter of the Earth.

OpenStudy (anonymous):

You just did this with a hamburger.....

OpenStudy (anonymous):

No, a pizza.

OpenStudy (paxpolaris):

\[\Large C= \pi \times d\] d=C/pi

OpenStudy (anonymous):

willie likes me tehe <3

OpenStudy (anonymous):

ok so it would be 7961?

OpenStudy (anonymous):

7962

OpenStudy (anonymous):

ok but my answers dont include any of those

OpenStudy (anonymous):

A. 7,957.8 miles B. 7,958.7 miles C. 7,958.6 miles D. 7,957.7 miles

OpenStudy (anonymous):

What they tell you to use for pi?

OpenStudy (anonymous):

7957.7 if you use the pi symbol on your calculator.

OpenStudy (anonymous):

circumfrance is 2pir=25000 r=25000/2pi r=3978.87 diameter d=2r=2*(3978.87)=7957.74

OpenStudy (anonymous):

@sami-21 You lose a bit of accuracy with that procedure because you are prematurely rounding off...